# Mixed Signals: A Diverse Point Cloud Dataset for Heterogeneous LiDAR V2X Collaboration

### v1

@@ 3.4 Dataset Analysis
Fig. 6 shows LiDAR intensity distributions from three sensors: RSU TOP, DOME, and Laser car sensors. DOME and TOP sensors record higher intensities because there is a large number of static objects (e.g., buildings, traffic lights) that are near them. In contrast, the Laser car sensor presents a smoother decline in intensity values because of its location on the vehicle, which allows the detection of objects at greater distances. EV-1 and EV-2 sensors do not capture intensity readings.
In our benchmark, we group 10 categories into 3 detection classes according to Table 2. Figure 8 shows the distribution of annotations of three classes with respect to their polar coordinate in the coordinate system of TOP. Figure 9 shows the distribution of dimensions and yaw angle of annotations of three classes. Figure 7 shows the number of annotations of each class in the training set and test set.
Figure 10 analyzes track lengths in the training and test set. For both splits, most tracks are under 10 seconds. This is due to the dynamic and typical speeds at the intersection environment. A sharp peak at 30 seconds indicates the presence of static objects detected primarily by the RSU for the entire sequence duration.
Figure 11 depicts the aggregation of point clouds from 5 agents and ground truth annotations in the coordinate system of TOP during a 4-second time span, which amounts to 40 time steps. The consistent pose of static objects and the smooth trajectory of dynamic objects visually demonstrate the quality of our annotation.
@@
Figure 7: Number of objects by class. The y-axis is displayed in log scale.
@@ 4 Proposed Tasks and Benchmarks
The inclusion of multiple agents and annotations in our dataset—in the form of 3D bounding boxes with track IDs enables the development of methods for various collaborative perception tasks, such as object detection, tracking, and motion forecasting. Given the importance of object detection in autonomous driving, we focus on collaborative detection tasks in this paper.
4.1 Definition of Tasks
Collaborative Object Detection. This is the classical collaborative object detection task [36, 22] where every connected agent (i.e., vehicles and RSUs) uses a shared model to (i) extract features from their point clouds, (ii) generate messages to send to other agents, and (iii) fuse the features of their point clouds with messages received from others. The goal is to detect every visible object in a region of interest. We define visibility by comparing the number of LiDAR points contained within an object’s bounding box to a threshold. In this task, these LiDAR points are sourced from any agents present within the region of interest.
Object Detection Enhanced by Communication to RSU. This task assumes that the RSU model is designed and trained by a different provider than the one responsible for the CAVs’ models. In this task, the RSU model is pre-trained in the single-vehicle detection setting to detect objects visible to its LiDARs. After the pre-training process, the RSU model is fixed. CAVs in the proximity of the RSU receive messages from the RSU to enhance their detection capabilities. The objective of this task is to detect all objects in a region of interest that are visible to either the CAV or the RSU.
@@ 4.1 Definition of Tasks
Table 3: Benchmarking results for the Collaborative Object Detection task. All fusion methods outperform the No Fusion baseline, highlighting the advantage of collaborative perception. Each fusion method involves trade-offs between detection performance and communication bandwidth overhead.
@@ 4.2 Benchmark
Evaluation Settings. Since the annotations are made in the coordinate system of TOP, we define the region of interest for the two detection tasks as the range $\left[-51.2,51.2\right]$ meters along both the x and y axes of this coordinate system. For evaluation, we transform objects detected by each agent into this coordinate system. The visibility threshold is set to 5 points. Since timestamp mismatches and localization errors are inherent in real-world applications and consequently present in our dataset, we do not artificially introduce them into the message exchanged among connected agents, as is often done in synthetic datasets [41, 40].
We measure object detection performance using Average Precision (AP). Detected objects are matched with ground truth based on their Intersection over Union (IoU) in the bird’s-eye view plane. A detection and a ground truth object are considered a match if their IoU exceeds thresholds of 0.3, 0.5, or 0.7.
In addition to AP, we measure the bandwidth consumption of each collaborative method to gauge their practicality. The total bandwidth consumption is calculated by multiplying the number of agents in the collaboration network by the size of the message each agent sends. While the number of agents is not dependent on the collaboration method of choice, the message size is. Therefore, we compute the bandwidth consumption by averaging the size of the messages that agents send, measured in Megabytes (MB). While some intermediate collaboration methods [36, 22, 40] employ specialized compressing algorithms to reduce the message size, other methods [41, 25, 7] do not. To obtain a fair comparison, we measure the size of uncompressed messages.
Methods. Our benchmark covers three conventional collaboration frameworks, namely Early fusion, Intermediate fusion [36, 41, 6, 40, 21, 14], and Late fusion, and the recent Laly fusion [7]. We detail implementation specifics in the Appendix of this report.
@@ 4.3 Results
We show the benchmark of Collaborative Object Detection task in Table 3. The result in this table clearly demonstrates the advantage of collaboration perception over single-agent perception as all fusion methods largely outperform No Fusion on every class. The comparison of three conventional fusion methods, including Early, Intermediate, and Late, shows that a higher precision is attained at the cost of a larger bandwidth consumption. In contrast, Laly fusion achieves comparable precision on Bike and Pedestrian compared to Early Fusion and Intermediate Fusion while consuming an order magnitude less bandwidth. The ability to achieve high performance for less bandwidth of Laly fusion coupled with its simplicity make this method a strong candidate for real-world deployment. However, we note that state-of-the-art methods still exhibit a performance gap on our dataset, suggesting the need for future algorithm design.
Table 4 presents the performance of different fusion methods on Object Detection Enhanced by Communication to RSU task. In this setting, detector training is more challenging, as each vehicle-centric detector must adapt to a frozen RSU detector. Nevertheless, results show that communication with RSU is still advantageous, as evidenced by the substantial performance improvement over the No Fusion baselines. Furthermore, the performance of the Laser car is better than the performance of the two EVs. This is because the LiDAR of the Laser car has a 360-degree coverage of its surroundings. On the other hand, the tilted angle of the LiDAR on the two EVs makes the region behind them unobservable. Additionally, the LiDARs on the two EVs do not capture the intensity information, resulting in a domain gap between their features and those from the RSU. These observations point to future research directions for developing methods that could work well with diverse sensor configurations.
@@ 5 Discussion and Conclusion
Our work presents the Mixed Signals V2X dataset, created through careful data selection, sensor synchronization and localization, and a strong investment in high quality annotations. To the best of our knowledge, our dataset is the first to support a heterogeneous sensor suite in an out-of-domain left hand side driving country, Australia, providing a diverse dataset addition to the field. We hope that the release of our dataset will facilitate research into complex and realistic settings for V2X perception. Future directions of research include studying communication protocols that ensure both fast transmission and directed communication that targets salient information.
@@ Acknowledgement
This research is funded by the University of Sydney – Cornell University Ignition Grants/Global Strategic Collaboration Awards, National Science Foundation NSF (IIS-2107161), and the New York Presbyterian Hospital. Minh-Quan Dao is funded by the ANNAPOLIS project managed by the French National Agency for Research (ANR-21-CE22-0014). Katie Luo is funded by the AAUW American Dissertation Fellowship.
[…]
In the appendix, we include extra details about the dataset and provided devkit, as well as annotation details and instructions given to annotators. We also provide additional sensor details.
@@
Table S1: Definitions of the annotation classes.
Appendix A Data and Devkit
@@ A.1 Devkit Description
We integrate our dataset into the framework OpenCOOD [41], which offers the implementation of various state-of-the-art collaborative perception methods. As OpenCOOD only provides single-class models, we adapt its implementation of Early, Intermediate, and Late Fusion models to detect three classes, including vehicles, bikes, and pedestrians. We added detection heads of 1-by-1 convolution layers to existing architectures to achieve this. In addition, we add the recent Laly fusion [7] to this framework. Every model in our benchmark uses PointPillar [20] as the backbone. Interested readers can refer to our code release for further details on architectures and training settings.
@@ A.2 Sample Data
Figure S1 shows an example of the collected data, where the points are colour-coded to represent the different LiDARs.
@@
Figure S1: Top-down view of the data collected at the location. LiDAR point clouds are colored by the vehicle and RSU that collected them, consisting of the 3 vehicle agents (red, yellow, and purple) and the Top and Dome LiDAR sensors of the RSU (green, blue). Best viewed in colour.
@@
The dataset aims to replicate realistic urban scenarios that reflect the complexities of real-world implementations by using multiple vehicles with diverse sensor configurations and a roadside unit. Real-world deployments of autonomous vehicles on streets incorporate LiDARs, which are becoming more affordable. Roadside infrastructure, such as roadside units, is also gaining popularity for traffic monitoring and data analytics, now often equipped with LiDAR, traffic light timing information, and communication systems to enhance robustness and applicability.
Our dataset consists of LiDAR point clouds, which offer the advantage of not capturing identifiable information like faces or license plates, thus preserving data privacy. This contrasts with camera images, which often require post-processing to anonymize sensitive details, potentially affecting data quality. Our dataset includes tracking IDs for each bounding box, and this information will be released alongside this paper. Benchmarks will be made available at a later date.
Appendix B Annotation Instructions
@@
We provide the instructions given to the Segments.ai***https://segments.ai/ annotators on the dataset website under annotation_instructions.pdf. We selected to invest in the quality of the annotations, applying rigorous quality control measures to guarantee accurate and consistent labeled data, minimizing errors, and maintaining high standards. In Table S1, we provide the definitions of the 10 fine-grained annotation classes in the Mixed Signals dataset. The breakdown of the fine-grain classes into the benchmarked classes can be found in the main text.
Appendix C Sensor Details
@@ C.1 Hardware and Synchronization Details
Table S2: Hardware specifications
The sensors in our multi-agent system were timestamped using GPS time as a common reference, and sensor details are provided in Table S2. The maximum time gap for matching sensor readings between 10 Hz rotational sensors is 50 ms. Since sensors rotate fully in 100 ms, angular positions differ by at most 180 degrees. If the time difference between readings were larger than 50 ms, it would be matched with the next or previous rotation instead. As shown in the original manuscript, precise sensor synchronization, robust multi-agent localization, and clearly defined annotation protocols produced high-quality data association across all sensors.
C.2 Definition of Heterogeneity in Sensor Suite

### v2

@@ 3.4 Dataset Analysis
Statistics. In our benchmark, we group 10 categories into 3 detection classes according to Table 2. Figure 7 shows the distribution of annotations of three classes with respect to their polar coordinate in the coordinate system of TOP. Figure 8 shows the distribution of dimensions and yaw angle of annotations of three classes. Figure 6 shows the number of annotations of each class in the training set and test set. Figure 9 analyzes track lengths in the training and test set. For both splits, most tracks are under 10 seconds. This is due to the dynamic and typical speeds at the intersection environment. A sharp peak at 30 seconds indicates the presence of static objects detected primarily by the RSU for the entire sequence duration. Figure 10 depicts the aggregation of point clouds from 5 agents and ground truth annotations in the coordinate system of TOP during a 4-second time span, which amounts to 40 time steps. The consistent pose of static objects and the smooth trajectory of dynamic objects visually demonstrate the quality of our annotation.
@@
Figure 6: Number of objects by class. The y-axis is in log scale.
@@ 4 Proposed Tasks and Benchmarks
Table 3: Benchmarking results for the Collaborative Object Detection task. All fusion methods outperform the No Fusion baseline, highlighting the advantage of collaborative perception. Each fusion method involves trade-offs between detection performance and communication bandwidth overhead. Models adapted from a premier R.H.S. V2V dataset [43] are shown in gray.
@@
Our dataset includes multiple agents and annotations in the form of 3D bounding boxes with track IDs. This enables the development of methods for various collaborative perception tasks, such as object detection, tracking, and motion forecasting. Given the importance of object detection in autonomous driving, we focus on collaborative detection tasks in the main text and report preliminary tracking benchmark results in Appendix B.2.
4.1 Definition of Tasks
Collaborative Object Detection. This is the classical collaborative object detection task [36, 22], where every connected agent (i.e., vehicles and RSUs) uses a shared model to (i) extract features from their point clouds, (ii) generate messages to send to other agents, and (iii) fuse the features of their point clouds with messages received from others. The goal is to detect every visible object in a region of interest. We define visibility by comparing the number of LiDAR points contained within an object’s bounding box to a threshold. In this task, these LiDAR points are sourced from any agents present within the region of interest.
Object Detection Enhanced by Communication to RSU. This task assumes that the RSU model is designed and trained by a different provider than the one responsible for the CAVs’ models. In this task, the RSU model is pre-trained in the single-vehicle detection setting to detect objects visible to its LiDARs. After the pre-training process, the RSU model is fixed. CAVs in the proximity of the RSU receive messages from the RSU to enhance their detection capabilities. The objective is to detect all objects in a region of interest that are visible to either the CAV or the RSU.
@@ 4.2 Benchmark
Evaluation Settings. Since the annotations are made in the coordinate system of TOP, we define the region of interest for the two detection tasks as the range $\left[-51.2,51.2\right]$ meters along both the x and y axes of this coordinate system. For evaluation, we transform objects detected by each agent into this coordinate system. The visibility threshold is set to 5 points. Since timestamp mismatches and localization errors are inherent in real-world applications and consequently present in our dataset, we do not artificially introduce them into the messages exchanged among connected agents (something that is often done in synthetic datasets [42, 41]). We measure object detection performance using Average Precision (AP). Detected objects are matched with ground truth based on their Intersection over Union (IoU) in the bird’s-eye view plane. A detection and a ground truth object are considered a match if their IoU exceeds thresholds of 0.3, 0.5, or 0.7. In addition to AP, we measure the bandwidth consumption of each collaborative method to gauge their practicality. The total bandwidth consumption is calculated by multiplying the number of agents in the collaboration network by the size of the message each agent sends. While the number of agents is not dependent on the collaboration method of choice, the message size is. Therefore, we compute the bandwidth consumption by averaging the size of the messages that agents send, measured in Megabytes (MB). While some intermediate collaboration methods [36, 22, 41] employ specialized compressing algorithms to reduce the message size, other methods [42, 25, 7] do not. To obtain a fair comparison, we measure the size of uncompressed messages.
Methods. Our benchmark covers three conventional collaboration frameworks, namely Early fusion, Intermediate fusion [36, 42, 6, 41, 21, 14], and Late fusion, and the recent Laly fusion [7]. We detail the benchmarking methodology specifics in the appendix.
@@ 4.3 Results
4.3.1 Collaborative Object Detection
We show the benchmark of the Collaborative Object Detection task in Table 3. The results in this table clearly demonstrate the advantage of collaboration perception over single-agent perception, as all fusion methods largely outperform No Fusion on every class. The comparison of three conventional fusion methods, including Early, Intermediate, and Late, shows that a higher precision is attained at the cost of a larger bandwidth consumption. In contrast, Laly fusion achieves comparable precision on Bike and Pedestrian compared to Early Fusion and Intermediate Fusion while consuming an order magnitude less bandwidth. The high performance at less bandwidth of Laly fusion coupled with its simplicity make this method a strong candidate for real-world deployment. However, we note that there is still ample room for improvement, particularly among the VRUs, suggesting the need for future algorithm design.
Domain Gap from prior R.H.T. Datasets. To illustrate the domain gap covered by Mixed Signals, we directly adapt an early-fusion and a late-fusion model trained on the right-hand traffic (R.H.T.) V2X dataset, V2V4Real [43], onto our dataset (grayed-out rows in Table 3). Performance degraded significantly, with vehicle headings often predicted incorrectly, indicating a learned prior from traffic flow (Figure 11). This highlights a substantial domain gap due to left-hand traffic and sensor modality differences, underscoring Mixed Signals’ unique contribution to the V2X perception landscape.
Figure 11: Detection visualization from adapting models trained on V2V4Real [43] into Mixed Signals. Ground truth bounding boxes are shown in green and predicted detections in blue, with heading indicated by the triangle. Observe that predicted heading directions are often aligned to priors learned in R.H.T. driving.
4.3.2 Detection via Communication to RSU
Table 4 presents the performance of different fusion methods on Object Detection Enhanced by Communication to RSU task. In this setting, detector training is more challenging, as each vehicle-centric detector must adapt to a frozen RSU detector. Nevertheless, results show that communication with RSU is still advantageous, as evidenced by the substantial performance improvement over the No Fusion baselines. Furthermore, the performance of the Laser car is better than the performance of the two EVs. This is because the LiDAR of the Laser car has a 360-degree coverage of its surroundings. On the other hand, the tilted angle of the LiDAR on the two EVs makes the region behind them unobservable. The LiDARs on the two EVs do not capture the intensity information, resulting in a domain gap between their features and those from the RSU. These observations point to future research directions for developing methods that could work well with diverse sensor configurations.
@@ 5 Discussion and Conclusion
Our work presents the Mixed Signals V2X dataset, created through careful data selection, sensor synchronization and localization, and a strong investment in high quality annotations. To the best of our knowledge, our dataset is the first to support heterogeneous sensor configurations with varying positions and orientations, collected in an out-of-domain left-hand traffic country, Australia, providing a diverse dataset addition to the field. We hope that the release of our dataset will facilitate research into complex and realistic settings for V2X perception. Future directions of research include studying communication protocols that ensure both fast transmission and directed communication that targets salient information. More broadly, we hope that our work could potentially further the research in the emerging, yet impactful, field of V2X communication.
@@ Acknowledgement
We thank Runsheng (Derrick) Xu and Hao Xiang for their insightful discussions and moral support throughout the early stage of this project. This research is funded by University of Sydney – Cornell University Ignition Grants/Global Strategic Collaboration Awards, National Science Foundation (IIS-2107161), and the New York Presbyterian Hospital. Minh-Quan Dao is funded by ANNAPOLIS project managed by the French National Agency for Research (ANR-21-CE22-0014), and Katie Luo by AAUW American Dissertation Fellowship.
Appendix
In this appendix material, we include: 1) extra details about the Mixed Signals dataset and the provided code devkit, 2) annotation details and instructions given to annotators, and 3) additional sensor details. We include an additional dataset teaser video in the dataset website***https://sites.coecis.cornell.edu/mixedsignals/#introvid that we encourage readers to watch.
@@
Table A1: Definitions of the annotation classes.
A Data and Devkit
@@ A.1 Devkit Description
We provide a separate devkit and additionally integrate our dataset into the framework OpenCOOD [42], which offers the implementation of various state-of-the-art collaborative perception methods. As OpenCOOD only provides single-class models, we adapt its implementation of Early, Intermediate, and Late Fusion models to detect three classes, including vehicles, bikes, and pedestrians. We added detection heads of 1-by-1 convolution layers to existing architectures to achieve this. In addition, we add the recent Laly fusion [7] to this framework. Every model in our benchmark uses PointPillar [20] as the backbone. Interested readers can refer to our devkit code releasehttps://github.com/chinitaberrio/Mixed-Signals and extended OpenCOOD integration for further details on architectures and training settings.
@@ A.2 Sample Data
Figure A1 shows an example of the collected data, where the points are colour-coded to represent the different LiDARs.
@@
Figure A1: Top-down view of the data collected at the location. LiDAR point clouds are colored by the vehicle and RSU that collected them, consisting of the 3 vehicle agents (red, yellow, and purple) and the Top and Dome LiDAR sensors of the RSU (green, blue). Best viewed in colour.
@@
The dataset aims to replicate realistic urban scenarios that reflect the complexities of real-world implementations by using multiple vehicles with diverse sensor configurations and a roadside unit. Real-world deployments of autonomous vehicles on streets incorporate LiDARs, which are becoming more affordable. Roadside infrastructure, such as roadside units, is also gaining popularity for traffic monitoring and data analytics, now often equipped with LiDAR, traffic light timing information, and communication systems to enhance robustness and applicability. Our dataset consists of LiDAR point clouds, which offer the advantage of not capturing identifiable information like faces or license plates, thus preserving data privacy. This contrasts with camera images, which often require post-processing to anonymize sensitive details, potentially affecting data quality. Our dataset includes tracking IDs for each bounding box, and this information will be released alongside this paper. Benchmarks will be made available at a later date.
Intensity Distributions. Figure A2 shows LiDAR intensity distributions from RSU TOP, DOME, and Laser car sensors. DOME and TOP sensors record higher intensities because there is a large number of static objects (e.g., buildings, traffic lights) near them. In contrast, the Laser car sensor presents a smoother decline in intensity values because of its location on the vehicle, which allows the detection of objects at greater distances. EV-1 and EV-2 sensors do not capture intensity readings. Therefore, a uniform approach to utilizing intensity values across all agent models would be inadequate.
Figure A2: Distribution of LiDAR intensities from RSU TOP, DOME, and Laser car sensors. Each sensor shows different intensity ranges and distributions. EV-1 and EV-2 LiDAR sensors do not have intensity readings.
B Annotation Instructions
@@
We provide the instructions given to the Segments.aihttps://segments.ai/ annotators in the attached material, titled “Spec Document - Multi-sensor labeling” at the bottom of the appendix. We selected to invest in the quality of the annotations, applying rigorous quality control measures to guarantee accurate and consistent labeled data, minimizing errors, and maintaining high standards.
B.1 Definitions of the Annotation Classes
The Mixed Signals dataset categories road agents in different vehicle types and pedestrians. Categories such as “Car” and “Truck” encompass common passenger and large transport vehicles, while “Emergency Vehicle” covers ambulances, fire trucks, and police cars, highlighting their importance in urban scenarios. “Bus” labels are designated for large passenger vehicles typically used in public transportation. The dataset also distinguishes between “Motorcycle” and “Motorized Bike,” and “Portable Personal Mobility Vehicle,” which includes modern personal transport devices like electric scooters and hoverboards. Traditional “Bicycle” labels account for both standard and electric bikes. Labels for “Electric Vehicle” and “Trailer” ensure that smaller, often data-collection vehicles and towable units are accurately represented. Finally, we labeled humans as “Pedestrians”. In Table A1, we provide the definitions of the 10 fine-grained annotation classes in the Mixed Signals dataset. The breakdown of the fine-grain classes into the benchmarked classes can be found in the main text.
B.2 Track Annotations and Multi-agent Tracking
We benchmark the performance of the planned tracking task for our dataset. The Mixed Signals dataset has labels for track ID’s, as seen in Figure 10 of the main text. We hope to include and benchmark tracking methods as an additional task which is supported by our dataset. We report some initial benchmarking results on the AB3DMOT tracking method [37] in Table A2. For further details about track labels, please explore the data itself; a distribution of the tracks are in Figure 9 of the main text.
Category sAMOTA AMOTA AMOTP
Vehicle 89.6 43.1 63.3
Pedestrian 76.6 32.6 42.8
Table A2: Tracking performance for AB3DMOT with V2X-ViT detections on the Mixed Signals validation split.
B.3 Annotation Details
The annotation process for this multi-sensor dataset involves handling joint scenes and synchronization discrepancies between sensors. Due to time synchronization, fast-moving objects might appear slightly offset across the data collected from different sensors. To address these discrepancies, annotators were instructed to prioritize the roadside unit point cloud for bounding box creation, following a set hierarchy. When there is a mismatch, bounding boxes should be aligned with the point cloud in the following order: roadside unit, EVs, and the urban vehicle. For example, if there is a difference between the roadside unit and the vehicles’ point cloud, the bounding box should only be fitted around the roadside unit points. This ensures consistency in object localization across frames despite synchronization lags.
C Sensor Details
@@ C.1 Hardware and Synchronization Details
Table A3: Hardware specifications.
Synchronization is especially important in dynamic environments, as any introduced time shifts can lead to positional inconsistencies, resulting in multiple detections of the same object. The sensors in our multi-agent system were timestamped using GPS time as a common reference, and sensor details are provided in Table A3. Rotational LiDARs continuously scan the environment in 360 degrees, thus, different portions of the surroundings are captured at slightly different times during a full rotation. When vehicles are in motion, their positions and orientations change dynamically between LiDARs sweeps. The maximum time gap for matching sensor readings between 10 Hz rotational sensors is 50 ms. Since sensors rotate fully in 100 ms, angular positions differ by at most 180 degrees. If the time difference between readings were larger than 50 ms, it would be matched with the next or previous rotation instead. As shown in the original manuscript, precise sensor synchronization, robust multi-agent localization, and clearly defined annotation protocols produced high-quality data association across all sensors.
C.2 Localization
Localization is one of the most critical tasks for CAV, estimating their position relative to a global reference frame. One of the most commonly used sensors for localization is the Global Navigation Satellite System (GNSS). GNSS offers access to a satellite constellation that provides global positioning via triangulation. However, despite its widespread use, GNSS has several drawbacks, particularly in urban environments. Its accuracy can be reduced in urban canyons, where tall buildings block or reflect signals, leading to degraded positioning accuracy. To overcome this problem, we use dense and accurate point cloud maps [31] as references for our localization algorithm.
C.3 Definition of Heterogeneity in Sensor Suite